OpenStudy (anonymous):

Please Please help me pass my math class? :) (multiple answers) Check all which apply: A radical is in simplest form when Choose at least one answer. a. There are no perfect squares contained in the radical. b. There are no radicals in the denominator of a fraction. c. You have estimated it to the nearest tenth. d. You have squared it.

OpenStudy (reemii):

it's a. ——————— explanation. here's a radical: \(\sqrt{Z}\) (coz if there's a perfect square inside, that is, if \(Z=n^2W\), then \(\sqrt{Z} = \sqrt{n^2W} = |n|\sqrt{W}\). something can "come out" of the sq.root

OpenStudy (primeralph):

I like girls.

OpenStudy (reemii):

as everybody!

OpenStudy (anonymous):

I dont like girls... lol. Is a the only answer? I can like pick as many as i want

OpenStudy (reemii):

ah, true. it's (a) and (b). coz: ppl don't let a fraction in this form: \(\frac{1}{\sqrt2}\), they write \(\frac{\sqrt2}{2}\) instead.

OpenStudy (anonymous):

Thanks!!

OpenStudy (primeralph):

@kika_rayne little advice, Once you got this, delete your question ASAP or you will be suspended for posting exam-related questions.

OpenStudy (anonymous):

Whaaat??

OpenStudy (primeralph):

Yeah. Six weeks at least.

OpenStudy (anonymous):

But its not for an exam.

OpenStudy (primeralph):

the title says "help me pass"

OpenStudy (reemii):

maybe it's "help me understand so I can pass in a few weeks" ? :°)

OpenStudy (anonymous):

haha im confused. but yeah i will close it